### v4

# Recognition of chordal graphs and cographs which are Cover-Incomparability graphs

Abstract
Cover-Incomparability graphs (C-I graphs) are an interesting class of graphs from posets. A C-I graph is a graph from a poset $P=(V,\leq)$ with vertex set $V$, and the edge-set is the union of edge sets of the cover graph and the incomparability graph of the poset. The recognition of the C-I graphs is known to be NP-complete (Maxová et al., Order 26(3), 229–236(2009)). In this paper, we prove that chordal graphs having at most two independent simplicial vertices are exactly the chordal graphs which are also C-I graphs. A similar result is obtained for cographs as well. Using the structural results of these graphs, we derive linear-time recognition algorithms for chordal graphs and cographs which are C-I graphs.
keywords:
Cover-Incomparability graphs; Chordal graph; Cographs; Recognition Algorithm.
1 Introduction
The cover-incomparability graphs of posets, or shortly C-I graphs form an interesting class of graphs from posets. These graphs were introduced in [8] as the underlying graphs of the so-called standard transit function of posets. The C-I graphs are precisely the graphs whose edge set is the union of edge sets of the cover graph and the incomparability graph (complement of a comparability graph) of a poset. Graph-theoretic characterization of C-I graphs has not been known until now and, moreover, the recognition complexity of C-I graphs is NP-complete (Maxová et al. [24]). Hence, problems in C-I graphs are focused on identifying the structure and characterization of well-known graph families, which are C-I graphs. Such C-I graphs studied include the family of split graphs, block graphs [7], cographs [9], Ptolemaic graphs [25], distance-hereditary graphs [25], and $k$-trees [23]. The C-I graphs were identified among the planar and chordal graphs along with new characterizations of the Ptolemaic graphs, respectively, in [4] and [3]. It is also interesting to note that every C-I graph has a Ptolemaic C-I graph as a spanning subgraph [4]. C-I graphs are also studied among the comparability graphs [1]. The effect of composition operation, lexicographic, and strong products of C-I graphs was studied in a recent paper [2].
[…]
Chordal graphs or triangulated graphs are graphs that do not have induced cycles of length greater than three. Together with cographs, which are exactly $P_{4}$-free graphs, they form two well-studied graph classes that have wider applications beyond mathematics. These two classes of graphs and their subclasses are also studied in many contexts, including theoretical, practical, and algorithmic interests. A vertex $v$ of a graph $G$ is a simplicial vertex if $v$ together with all its adjacent vertices forms a clique (a complete subgraph) in $G$. It is a well-known fact that every chordal graph contains at least one simplicial vertex, and a non complete chordal graph contains at least two nonadjacent simplicial vertices. Chordal graphs that contain exactly two non-adjacent simplicial vertices form a special class of chordal graphs.
The recognition algorithm for an arbitrary chordal graph [30, 27] and a cograph [14] is well known and has linear-time complexity. In this paper, we identify chordal graphs having exactly two non-adjacent simplicial vertices and prove that they are C-I graphs. In other words, we prove that the intersection of the class of chordal graphs and the class of C-I graphs is precisely the class of chordal graphs having exactly two non-adjacent simplicial vertices. We further develop a linear-time algorithm for recognizing such chordal graphs. It is interesting to observe, from [4, 7], that in the class of C-I graphs, the classes of chordal, strongly chordal, and interval graphs are one and the same.
In this paper, similar to the chordal C-I graphs, we develop a linear-time recognition algorithm for cographs which are also C-I graphs from the structural results already proved in [9]. For this algorithm, we make use of the special type of cotree structure of cographs.
Chordal graphs are precisely the intersection graphs of subtrees of a tree [20]. Most of the information contained in a chordal graph is captured in its clique tree representation, which is useful for its algorithmic applications [20, 29].
Cographs are exactly the $P_{4}$-free graphs and the class of cographs has been intensively studied since its definition by Seinsche [28]. The cographs appear as comparability graphs of series-parallel partial orders [22], and can be generated from the single-vertex graph $K_{1}$ by complementation and disjoint union operations. It is well known that any cograph has a canonical tree representation called a cotree. This tree decomposition scheme for cographs is a particular case of modular decomposition [19] that applies to arbitrary graphs. Indeed, the algorithm that computes the modular tree decomposition of an arbitrary graph in linear-time can also recognize cographs in linear-time. In 1994, linear-time modular decomposition algorithms were designed independently by Cournier and Habib [15] and by McConnell and Spinrad [26]. In 2001, Dahlhaus et al. [16] proposed a simpler algorithm. In 2004, Habib and Paul [21] proposed a new algorithm, which is not incremental, and instead of building the cotree directly, it first computes a special ordering of the vertices, namely a factorizing permutation, using a very efficient partition refinement techniques via two elementary refinement rules.
A search of a graph visits all vertices and edges of the graph and will visit a new vertex only if it is adjacent to some previously visited vertex. The two fundamental search strategies are Breadth-First Search (BFS) and Depth-First Search (DFS). As the names indicate, BFS visits all previously unvisited neighbors of the currently visited vertex before visiting the previously unvisited non-neighbors. Several greedy recognition algorithms for chordal graphs are known. The most famous is Lex-BFS [27], a variant of $BFS$, introduced by Rose et al. in [27] and Maximum Cardinality Search (MCS for short) [30]. Both algorithms are linear. The recognition of chordal graphs involves two distinct phases: the execution of MCS or Lex-BFS in order to compute an elimination ordering and a checking procedure to decide whether this elimination ordering is perfect (PEO). By employing Lex-BFS as a basic method to check the chordality and to find the perfect elimination ordering, the recognition algorithm for the chordal graphs can be done in linear-time for chordal graphs which are C-I graphs. For cographs, using the BFS, we can check whether a given rooted tree is the cotree of the C-I cograph in linear-time because of the special structure of the cotrees of these graphs.
The rest of this section is organized as follows. In Section  2, we begin with some preliminaries. In Section 3, we characterize chordal C-I graphs in terms of the number of independent simplicial vertices and present a linear-time algorithm for recognizing chordal C-I graphs based on this characterization. In Section 4, we present a structure of the C-I cograph and characterize the structure of the cotree of C-I cographs and, using this, present a linear-time recognition algorithm for C-I cographs.
2 Preliminaries
A partially ordered set or poset $P=(V,\leq)$ consists of a nonempty set $V$ and a reflexive, antisymmetric, transitive relation $\leq$ on $V$, denoted as $P=(V,\leq)$. We call $u\in V$ an element of $P$. If $u\leq v$ or $v\leq u$ in $P$, we say $u$ and $v$ are comparable, otherwise incomparable. If $u\leq v$ but $u\neq v$, then we write $u<v$. If $u$ and $v$ are in $V$, then $v$ covers $u$ in $P$ if $u<v$ and there is no $w$ in $V$ with $u<w<v$, denoted by $u\lhd v$. We write $u\lhd\lhd v$ if $u<v$ but not $u\lhd v$. By $u||v$, we mean that $u$ and $v$ are incomparable elements of $P$. In this paper, we consider only posets defined on finite sets. Let $V^{\prime}\subseteq V$ and $Q=(V^{\prime},\leq^{\prime})$ be a poset, $Q$ is called a subposet of $P$, if $u\leq^{\prime}v$ if and only if $u\leq v$, for any $u,v\in V^{\prime}$. The subposet $Q=(V^{\prime},\leq)$ is a chain (antichain) in $P$, if every pair of elements of $V^{\prime}$ is comparable (incomparable) in $P$. A chain of maximum cardinality is named as the height of $P$ denoted as $h(P)$. An element $u$ in $P$ is a minimal (maximal) if there is no $x\in V$ such that $x\leq u(u\leq x)$ in $P$. A finite ranked poset (also known as graded poset [11]) is a poset $P=(V,\leq)$ that is equipped with a rank function $\rho:V\rightarrow\mathbb{Z}$ satisfying:
  • $\rho$ has value 00 on all minimal elements of $P$, and
  • $\rho$ preserves covering relations: if $a\lhd b$ then $\rho(b)=\rho(a)+1$.
A ranked poset $P$ is said to be complete if for every $i$, every element of rank $i$ covers all elements of rank $i-1$. For a completely ranked poset $P=(V,\leq)$ we say that the element $v\in V$ is at height $i$ if $\rho(v)=i-1$. For disjoint posets, $P_{1}=(V_{1},\leq)$ and $P_{2}=(V_{2},\leq)$, the sum of the posets, $Z=P_{1}+P_{2}$ is defined as the poset $(V_{1}\cup V_{2},\leq)$, where $z_{1}\in z_{2}$ in $Z$, if and only if $z_{1}\leq z_{2}$ in $P_{1}$ or $z_{1}\leq z_{2}$ in $P_{2}$. We refer to [11], for notions of posets.
Let $G=(V,E)$ be a connected graph, vertex set and edge set of $G$ denoted as $V(G)$ and $E(G)$ respectively, the complement of $G$ is denoted as $\overline{G}$. For a vertex $v\in V(G)$, the set of all vertices adjacent to $v$ is called the open neighborhood of $v$ and is denoted by $N(v)$. The set consisting of the open neighborhood and the vertex $v$ is the closed neighborhood of $v$ and is denoted by $N[v]$. A tree is a graph in which two vertices are connected by exactly one path. Let $T$ be a rooted tree and two vertices $x$ and $y$ in $T$, we say that $x$ is an ancestor of $y$ and $y$ is a descendant of $x$ if $x$ lies on the path from $y$ to the root of $T$. For a set of leaves $S$ of $T$, we say that the lowest common ancestor (LCA) of $S$ is the internal node $v$ of $T$ such that $v$ is the root of the smallest rooted subtree of $T$ containing $S$. A graph $H$ is said to be a subgraph of $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. $H$ is an induced subgraph of $G$ if for $u,v\in V(H)$ and $uv\in E(G)$ implies $uv\in E(H)$. A graph $G$ is said to be H-free if $G$ has no induced subgraph isomorphic to $H$. A complete graph is a graph whose vertices are pairwise adjacent, denoted $K_{n}$, a set $S\subseteq V(G)$ is a clique if the subgraph of $G$ induced by $S$ is a complete graph, and a maximum clique is a clique that is not contained by any other clique. A vertex $v$ is called simplicial vertex if its neighborhood induces a complete subgraph. An independent set in a graph is a set of pairwise non-adjacent vertices. If graphs $G_{1}$ and $G_{2}$ have disjoint vertex set $V_{1}$ and $V_{2}$ and edge set $E_{1}$ and $E_{2}$ respectively, then their union $G=G_{1}\cup G_{2}$ has $V=V_{1}\cup V_{2}$ and $E=E_{1}\cup E_{2}$ and their join, denoted by $G_{1}\vee G_{2}$, consists of $G=G_{1}\cup G_{2}$ and all edges joining $V_{1}$ with $V_{2}$.
A graph $G$ is chordal if it contains no induced cycles of length more than 3. A graph $G$ is distance-hereditary if every induced path is also the shortest path in $G$. A graph $G$ is Ptolemaic if it is distance-hereditary and chordal. Equivalently, $G$ is Ptolemaic if and only if it is a 3-fan-free chordal graph. $P_{4}$-free graphs are called cographs. A graph $G$ that is both chordal and cograph is called chordal cograph, also known as trivially perfect graph. Let $P=(V,\leq)$ be a poset. A graph $G$ is the cover graph of $P$ if $V(G)=V$ and $uv\in E(G)$ if and only if either $u\lhd v$ or $v\lhd u$ in $P$. Similarly, $G$ is the comparability graph of $P$ if $V(G)=V$ and $uv\in E(G)$ if and only if $u$ and $v$ are comparable in $P$. The incomparability graph is the complement of the comparability graph. Finally, the cover-incomparability graph (C-I graph) of a poset $P=(V,\leq)$ denoted as $G_{P}$ is the graph $G=(V,E)$, where $uv\in E(G)$, if $u\lhd v$ or $v\lhd u$ or $u||v$ in $P$. A graph is a C-I graph if it is the C-I graph of some poset $P$. It may be noted that the class of C-I graphs is not a hereditary class in the sense that every induced subgraph of a C-I graph need not be a C-I graph. We call a graph that is both a chordal and a C-I graph as a chordal C-I graph. Similarly, we use the term Ptolemaic C-I graph, C-I cograph, chordal C-I cograph, etc. to denote the C-I graph which is Ptolemaic, cograph, chordal cograph, etc.
Fig. 1: (a) A poset $P$, (b) incomparability graph of the poset $P$, and (c) cover-incomparability graph of the poset $P$.
In Figure 1 (a), we illustrates the Hasse diagram of a poset $P$, which is isomorphic to the cover graph of $P$. Figure 1(b), depicts the incomparability graph of the poset $P$, and in Figure 1(c), we depicts the C-I graph of $P$. As already mentioned, the class of C-I graphs is not hereditary; for example, the 4-fan is a C-I graph (as shown in Figure 1(c)), but the claw graph is an induced subgraph of this graph, which is not a C-I graph.
Now we recall some basic properties of posets and their C-I graphs.
Lemma 1.
[8] Let P be a poset. Then
  1. (i)
    the C-I graph of P is connected;
  2. (ii)
    points of P that are independent in the C-I graph of P lie on a common chain;
  3. (iii)
    an antichain of P corresponds to a complete subgraph in the C-I graph of P;
  4. (iv)
    the C-I graph of P contains no induced cycles of length greater than 4.
Lemma 2.
[25] If $G$ is a C-I graph, then $G$ does not contain 3 independent simplicial vertices.
Lemma 3.
[25] Let $P$ be a poset and $G$ its C-I graph. If $v$ is a simplicial vertex in $G$, then $v$ is a maximal or a minimal element of $P$.
Theorem 1.
[23] Let $G=(V,E)$ be a C-I graph of a poset $P$ and $v\in V$ a minimal or maximal element in $P$. Then $G\backslash v$ is also a C-I graph.
In the following sections, we discuss algorithms for recognizing chordal C-I graphs and C-I cographs.
3 Structure of chordal C-I graphs
In this section, we study the structure of chordal C-I graphs. For that, we prove that a chordal graph having exactly two independent vertices is a chordal graph that is also a C-I graph.
If $v$ is a simplicial vertex of a chordal graph $G$, then the closed neighborhood $N[v]$ is a clique in $G$. So removal of $v$ does not affect the chordal property of $G\setminus\{v\}$. This is stated in the following remark.
Remark 1.
If $G$ is a chordal graph and $v$ is a simplicial vertex, then $G\setminus v$ is chordal.
Lemma 4.
Let $G$ be a chordal graph, and $v$ be a simplicial vertex of $G$. If $G$ has exactly two independent simplicial vertices, then either $G\setminus v$ is chordal and has exactly two independent simplicial vertices or $G\setminus v$ is a complete graph.
Proof.
Let $G$ be a chordal graph with exactly two independent simplicial vertices and $v$ be a simplicial vertex. Then by Remark 1, $G\setminus v$ is a chordal graph. Since $v$ is a simplicial vertex, the removal of $v$ from $G$ does not increase the number of independent simplicial vertices in $G\setminus v$. And if $G\setminus v$ is not a complete graph, then removal of $v$ from $G$ does not decrease the number of independent simplicial vertex in $G\setminus v$. That is, $G\setminus v$ has exactly two independent simplicial vertices, otherwise $G\setminus v$ is a complete graph. Hence the lemma. ∎
From Lemmas 3-4 and Theorem 1, we get the following corollary.
Corollary 1.
Let $G$ be a chordal C-I graph and $v$ be a simplicial vertex of $G$. If $G$ has exactly two independent simplicial vertices, then $G\setminus v$ is a chordal C-I graph and has exactly two independent simplicial vertices or $G\setminus v$ is a complete graph.
A pendant clique of a graph $G$ is a clique that contains a clique separator $C$ such that one of the components obtained after removing $C$ is a single vertex. Observe that if $G$ is a chordal graph, then $G$ has a pendant clique.
Let $G$ be a chordal graph having at most two independent simplicial vertices. We denote by $\mathscr{G}$ the family of chordal graphs obtained from $G$ by adding a vertex $v$ to $G$ such that $v$ is adjacent to some or all vertices in a pendant clique of $G$.
The following remark is immediate.
Remark 2.
Let $G^{\prime}$ be a chordal graph having exactly two independent simplicial vertices. If $G$ is the graph obtained by adding a simplicial vertex $v$ to $G^{\prime}$ such that the resulting graph has exactly two independent simplicial vertices, then $G$ belongs to the family $\mathscr{G}$.
Lemma 5.
Let $G^{\prime}$ be a chordal C-I graph having exactly two independent simplicial vertices, and let $G$ be a graph obtained by adding a simplicial vertex $v$ to $G^{\prime}$. If $G$ has exactly two independent simplicial vertices, then $G$ is a chordal C-I graph.
Proof.
Let $G$ be a chordal graph obtained by adding a simplicial vertex $v$ to a chordal $G^{\prime}$ that has exactly two independent simplicial vertices such that $G$ also has exactly two independent simplicial vertices. Then by Remark 2, the vertex $v$ is such that $v$ is adjacent to some or all vertices in a pendant clique in $G^{\prime}$.
Now we need to prove that $G$ is a chordal C-I graph. Since $G^{\prime}$ is both a C-I graph and chordal, let $P^{\prime}$ be a poset such that $G_{P^{\prime}}\cong G^{\prime}$. Since the vertex $v$ is added to some or all vertices in a pendant clique in $G^{\prime}$, the graph $G$ is chordal. Since $G_{P}^{\prime}$ is chordal and has exactly two independent simplicial vertices, $G_{P}^{\prime}$ has exactly two pendant cliques, and the two pendent cliques are formed, respectively, by $S$ and $S^{\prime}$, where $S$ and $S^{\prime}$ are defined as follows.
$M=\{u\in P^{\prime}\mid u\text{ is a maximal element of }P^{\prime}\}$, $S_{1}=\{w\in P^{\prime}\mid w\lhd u,u\in M\}$, $M^{\prime}=\{u\in P^{\prime}\mid u\text{ is a minimal element of }P^{\prime}\}$, $S_{1}^{\prime}=\{w\in P^{\prime}\mid u\lhd w,u\in M^{\prime}\}$ $S=M\cup S_{1}\setminus\{u\in S_{1}\mid\exists w\in S_{1},u\lhd\lhd w\text{ in }P^{\prime}\}$ and
$S^{\prime}=M^{\prime}\cup S_{1}^{\prime}\setminus\{u\in S_{1}^{\prime}\mid\exists w\in S_{1}^{\prime},w\lhd\lhd u\text{ in }P^{\prime}\}$. Now $v$ is adjacent to some or all vertices of the set $S$ or $S^{\prime}$ in $G$. If $v$ is adjacent to some vertices of $S$, then $v$ must be adjacent to all the elements of $M$, as otherwise there will be more than two independent simplicial vertices. Similarly, if $v$ is adjacent to some vertices of $S^{\prime}$, then $v$ must be adjacent to all elements of $M^{\prime}$. Now, we construct a poset $P$ from $P^{\prime}$ as follows.
If $v$ adjacent to some vertices of $S$:
  • If $v$ is adjacent to only the elements in $M$, then $P$ is constructed from $P^{\prime}$ with the covering relation defined as $u\lhd v$ for all $u\in M$.
  • If $v$ is adjacent to the elements of $S_{0}$, where $M\subset S_{0}\subset S$, then $P$ is constructed from $P^{\prime}$ with the covering relation as $u^{\prime}\lhd v$ for all $u^{\prime}\in S_{0}\setminus M$ and $u^{\prime\prime}\lhd u\lhd v$ for all $u^{\prime\prime}\in S\setminus S_{0}$ for some $u\in M$.
  • If $v$ is adjacent to all elements of $S$, then $P$ is constructed from $P^{\prime}$ with the covering relation defined as $u\lhd v$ for all $u\in S\setminus M$.
Similarly, if $v$ adjacent to some vertices of $S^{\prime}$:
  • if $v$ is adjacent to only the elements in $M^{\prime}$ then $P$ is constructed from $P^{\prime}$ with the covering relation defined as $v\lhd u$ for all $u\in M^{\prime}$.
  • if $v$ is adjacent to the elements of $S_{0}^{\prime}$, where $M^{\prime}\subset S_{0}^{\prime}\subset S^{\prime}$, then $P$ is constructed with the covering relation defined as $v\lhd u^{\prime}$ for all $u^{\prime}\in S_{0}^{\prime}\setminus M^{\prime}$ and $v\lhd u\lhd u^{\prime\prime}$ for all $u^{\prime\prime}\in S^{\prime}\setminus S_{0}^{\prime}$ and some $u\in M^{\prime}$.
  • If $v$ is adjacent to all elements of $S^{\prime}$, then $P$ is constructed from $P^{\prime}$ with the covering relation $v\lhd u$ for all $u\in S^{\prime}\setminus M^{\prime}$.
It follows from construction that $P$ is a well-defined poset and that $G$ is isomorphic to $G_{P}$. That is, $G$ is a chordal C-I graph. Hence the lemma. ∎
A perfect elimination ordering (PEO) is an ordering $\pi=v_{1},\ldots,v_{n}$ of vertices in $G$ such that the neighborhood $N[v_{i}]$ of $v_{i}$ is a clique of the subgraph $G_{\{v_{i},\ldots,v_{n}\}}$ induced by the vertices $\{v_{i},\ldots,v_{n}\}$ of $G$. The following characterizations of chordal graphs are well known.
Theorem 2.
[18] A graph $G$ is chordal if and only if $G$ has a perfect elimination ordering.
Lemma 6.
[17] Every chordal graph $G$ has a simplicial vertex. If $G$ is not complete, then it has two non-adjacent simplicial vertices.
In the following, we prove an interesting characterization of C-I chordal graphs as precisely those chordal graphs having exactly two independent simplicial vertices.
Theorem 3.
Let $G$ be a chordal graph. Then $G$ is a C-I graph if and only if $G$ is a complete graph or $G$ has exactly two independent simplicial vertices.
Proof.
Suppose $G$ is a chordal and C-I graph, then $G$ is a complete graph, or $G$ contains exactly two independent simplicial vertices by Lemma 2 and 6.
Conversely, every complete graph is a C-I graph. It remains to prove that if $G$ is chordal and $G$ has exactly two independent simplicial vertices, then $G$ is a C-I graph.
Let $G$ be the chordal graph with the vertex set $V(G)=\{v_{1},v_{2},v_{3},\ldots,v_{n}\}$. Since $G$ is chordal, there is a perfect elimination ordering(PEO), let $v_{n},v_{n-1},\ldots,v_{2},v_{1}$ be a PEO in $G$. By the definition of PEO on a chordal graph, the neighborhood $N(v_{n-i})$ of $v_{n-i}$ is a clique in the subgraph $G_{\{v_{n-\{i+1\}},\ldots,v_{1}\}}$, for $i=0,1,\ldots n-1$ and also the subgraphs $G_{\{v_{n-(i+1)},\ldots,v_{1}\}}$ are chordal for $i=0,1,\ldots n-1$.
According to Lemma 4, if a chordal graph has exactly two independent simplicial vertices, then the removal of a simplicial vertex results in either a chordal graph with exactly two independent simplicial vertices or a complete graph. We eliminate the simplicial vertices until we get the smallest non-trivial chordal graph containing exactly two independent simplicial vertices, say $G_{k}$. The removal of a simplicial vertex from $G_{k}$ results in a complete graph. That is, $G_{k-1}$ is a complete graph with $G_{k-1}=G_{k}\setminus{v_{k}}$, where $v_{k}$ is a simplicial vertex in $G_{k}$. Clearly, $G_{k-1}$ is a C-I graph being a complete graph. Now, we add the simplicial vertex $v_{k}$ to $G_{k-1}$ and form the graph $G_{K}$ by adding the deleted edges back. We continue this process by adding the eliminated simplicial vertices $v_{k+1},v_{k+2},\ldots v_{n}$ and deleted edges in the reverse order in the PEO to the graphs $G_{k},G_{k+1},\ldots G_{n-1}$, respectively obtaining finally the graph $G_{n}\cong G$.
It may be noted that we add edges so that the resulting graph obtained in each stage contains exactly two independent simplicial vertices.
In particular, the graph $G_{k}$ is obtained by adding the simplicial vertex $v_{k}$ to $G_{k-1}$. The vertex $v_{k}$ is adjacent to some vertices of $G_{k-1}$. Let $C_{k-1}\subset V(G_{k-1})$ and $C_{k-1}^{\prime}=V(G_{k-1})\setminus C_{k-1}$ such that $v_{k}$ is adjacent to every element of $C_{k-1}$ and not adjacent to any element of $C_{k-1}^{\prime}$. We form a poset, say $P_{0}$ consisting of elements, $C_{k-1}^{\prime}$, $v_{k}$, and $C_{k-1}$. Now fix $C_{k-1}^{\prime}$ as the set of minimal elements, $v_{k}$ covering every element in $C_{k-1}^{\prime}$ and all elements of $C_{k-1}$ covering $v_{k}$. It is clear that the C-I graph of $P_{0}$ is isomorphic to $G_{k}$. That is, we obtain that $G_{k}$ is a C-I graph. It is in-fact a Ptolemaic C-I graph of the completely ranked poset with rank 2, where $C_{k-1}^{\prime}$ represents elements of rank 0, $v_{k}$ represents the element of rank 1 and $C_{k-1}$ represents elements of rank 2. Thus $G_{k}$ is a chordal C-I graph with exactly two simplicial vertices.
Now the graph $G_{k+1}$ is obtained by adding the simplicial vertex $v_{k+1}$ and the deleted edges to $G_{k}$ and from Lemma 5, we obtain that $G_{k+1}$ is a chordal C-I graph having exactly two independent simplical vertices. By by the inductive arguments, we have that $G_{k+2},G_{k+3},\ldots,G_{n}$ are chordal C-I graphs with exactly two independent simplicial vertices. Finally, we get that the graph $G_{n}$ is isomorphic to $G$ and hence $G$ is a C-I graph. Hence the theorem.
3.1 Recognition of chordal C-I graphs
In this section, we design an algorithm for recognizing chordal C-I graphs. For this, we use the clique tree representation of a chordal graph. Since this algorithm computes a clique tree, if the graph is a chordal graph explicitly, it turns out that the algorithm provides a certification. It is well known that a chordal graph $G$ can be characterized as the intersection graph of subtrees of some tree $T$. Such a tree is known as the clique tree of $G$. This well-known theorem is proved independently by Buneman [12] and Gavril [20]. A tree representation of a chordal graph $G$ is a pair $(T,\mathscr{F})$ where $T$ is a tree and $\mathscr{F}$ is a family of subtrees of $T$ such that the intersection graph of $\mathscr{F}$ is isomorphic to $G$. Further Gavril [20] has shown that given a chordal graph $G$, it is possible to construct a tree $T$ with vertex set $K=\{q_{1},q_{2},\ldots,q_{r}\}$ where $q_{i}$ corresponds to the maximal clique $Q_{i}$ of $G$, such that $(T,\{R_{v_{1}},R_{v_{2}},\ldots,R_{v_{n}}\})$ is a tree representation of $G$. Here, each $R_{v_{i}}$, $1\leq i\leq n$, is the set of maximal cliques that contain the vertex $v_{i}$, that is, $R_{v_{i}}=\{q_{j}\mid v_{i}\in Q_{j}\}$. Such a tree representation of $G$ is called a clique tree of $G$. The clique tree $T$ of $G$ need not be unique, for the sets $R_{i}$ determine the edges of $T$, not necessarily in a unique manner. Fig.2 shows a chordal graph and two of its clique trees.
Fig. 2: A chordal graph and two of its clique trees
Theorem 4 (Gavril [20] and Buneman [12]).
The following propositions are equivalent:
  1. 1.
    $G$ is a chordal graph.
  2. 2.
    $G$ is the intersection graph of a family of subtrees of a tree.
  3. 3.
    There exists a tree $T$(called clique tree) with vertex set $\{q_{1},q_{2},\ldots,q_{r}\}$ such that for each vertex $v$, the set $R_{v}=\{q_{i}\mid v\in Q_{i}\}$ induces a subtree of $T$. Here $q_{i}$ represents the maximal clique $Q_{i}$.
From the above description, we have that a graph $G$ is a chordal graph if and only if there exists a clique tree $T$ with maximal cliques of $G$ as the vertices of $T$ and any two cliques containing $v\in V(G)$ are either adjacent in $T$ or connected by a path of cliques that contain $v$. The Lemma below is also very useful in understanding the structure of chordal graphs and their simplicial vertices.
Lemma 7.
[5] A vertex is simplicial if and only if it belongs to precisely one maximal clique.
Even though the result of the following lemma is a known property of chordal graphs, we require this for establishing the recognition of chordal graphs which are C-I graphs, and hence we state it below.
Lemma 8.
Every leaf node of a clique tree of a chordal graph contains a simplicial vertex.
Proof.
Let $T$ be a clique tree of a chordal graph $G$. Let $C_{k}$ be a leaf node of $T$. Since $C_{k}$ is a leaf node of $T$ there is a node $C_{k^{\prime}}$ in $T$ such that $C_{k}C_{k^{\prime}}\in E(T)$. Since $T$ is a clique tree of $G$ every node of $T$ is a maximal clique of $G$, and hence there is a vertex $v$ of $G$ in $C_{k}$ but not in $C_{k^{\prime}}$. From the definition of the clique tree, it is clear that $v$ is not in any other nodes of $T$ (cliques). Then it follows that $v$ is a simplicial vertex by Lemma 7. ∎
The following lemma will also be useful for the recognition of chordal C-I graphs.
Lemma 9.
Let $G$ be a chordal graph with the clique tree $T$ being a path. Let $C_{i}$ be an internal node of $T$, $C_{i+1}$, the parent node of $C_{i}$, and $C_{i-1}$ the child node of $C_{i}$ in $T$. If $v$ is vertex of $G$ such that $v\in C_{i}\setminus C_{i+1}$ and $v\in C_{i}\setminus C_{i-1}$, then $v$ is a simplicial vertex in $G$.
Proof.
Since $v\in C_{i}\setminus C_{i+1}$ and $v\in C_{i}\setminus C_{i-1}$ and $T$ is a path, it follows from the definition of a clique tree that $v$ is not in any other cliques in $T$. Then by Lemma 7, $v$ is a simplicial vertex. ∎
Based on Theorem  3 and Lemma 8 and 9, we formulate the following algorithm for recognizing a chordal C-I graph $G$. The correctness of the algorithm follows from these results.
Input : $G$ be a connected graph with $|V(G)|=n$ and $|E(G)|=m$
Output : $G$ is chordal C-I graph or not
  • 1Step 1:
    Apply the perfect elimination ordering (PEO) on $G$ and check whether $G$ is a chordal graph or not.
If there is no PEO then return $G$ not chordal and stop. Otherwise, go to Step 2.
  • 2Step 2:
    Using the PEO find the clique tree $T$ of the given graph $G$ using Blair-Peyton algorithm.
  • 3 If the clique tree has more than 2 leaf nodes, then return $G$, not chordal C-I graph, and stop.
    Otherwise, go to Step 3. (Now the clique tree is a path, and the vertices are in order $C_{1},C_{2},\ldots,C_{k}$)
  • 4Step 3:
    Check whether there is an element in $C_{i}$ that is not in $C_{i+1}$ and $C_{i-1}$ for $i=2,3,\ldots,k-1$.
  • 5 If such an element exists at any stage, then stop and return $G$ as not a chordal C-I graph.
    Otherwise, return $G$ as a chordal C-I graph.
    Algorithm 1 Algorithm for recognizing given graph $G$ is chordal C-I graph $G$ or not.
    The time complexity of Algorithm 1 can be analyzed as follows.
    Using Lex-BFS or MCS, a PEO can be found in $O(n+m)$ time. In Step 2, the Blair-Peyton algorithm is used to construct a clique tree from a chordal graph based on the PEO [5]. This algorithm also runs in linear time, $O(n+m)$. Additionally, checking whether there are more than two leaf nodes can be done in constant time.
    In Step 3, let $k$ be the length of the clique tree which is a path $P$. Then, in the worst case, the size of the sets $C_{i}$ is of the order $O(n/k)$. To check whether an element $v\in C_{i}\setminus C_{i+1}$ and $v\in C_{i}\setminus C_{i-1}$ takes $O(3n/k)$ time by using a hash set operation. Since there are at most $k-2$ internal nodes, the total time for Step 3 is $O(n)$. Now, the complexity of Algorithm 1 is $O(n+m)$.
    According to Theorem 3, a graph $G$ is a chordal C-I graph if and only if $G$ is either a complete graph or has exactly two independent simplicial vertices. Lemma 7 and Lemma 8 further imply that the clique tree of a chordal C-I graph contains exactly two leaf nodes, meaning that the clique tree forms a path. The next step is to certify that none of the internal vertices of the clique tree contain simplicial vertices. Lemma 7 and Lemma 9 address this in Step 3 of the algorithm.
    Since the Blair-Peyton algorithm explicitly computes the clique tree when $G$ is a chordal graph and after Steps 2 and 3 of the algorithm outputs a clique tree which is a path and contains no internal vertices with simplicial vertices, if the given graph is a chordal C-I graph. So the algorithm actually returns a chordal C-I graph if $G$ is such a graph, which is a certification model.
    4 Structure of C-I cographs
    In this section, we present an algorithm for recognizing C-I cograph.
    Theorem 5.
    [9] Let $G$ be a chordal cograph. Then $G$ is a C-I graph if and only if $G$ is a connected graph that contains at most two maximal cliques.
    From Theorem 5, we get the following remark
    Remark 3.
    A graph $G$ is a chordal C-I cograph if and only if there exists three pairwise disjoint sets $C_{1},C_{2}$ and $C_{3}$ such that $V(G)=C_{1}\cup C_{2}\cup C_{3}$, and $x,y\in V(G)$ are adjacent in $G$ if and only if $x,y\in C_{1}\cup C_{2}$ or $x,y\in C_{2}\cup C_{3}$.
    That is, $C_{1},C_{2},C_{3},C_{1}\cup C_{2}$ and $C_{2}\cup C_{3}$ form cliques and the graph has no other edges. So in the C-I chordal cograph, if $xy,xz\notin E(G)$ then $yz\in E(G)$. It is clear that a chordal C-I cograph is a Ptolemaic C-I graph as it is a C-I graph of a completely ranked poset of height 3. The structure of an arbitrary chordal C-I cograph is shown in Fig. 3.
    Fig. 3: General structure of a chordal C-I cograph $G$
    Theorem 6.
    [9] A graph $G$ is a C-I cograph if and only if $G$ is the join of chordal C-I cographs.
    From Theorem 6, we get the following remark.
    Remark 4.
    Let $G$ be a C-I cograph such that $G=G_{1}\vee G_{2}\vee\cdots\vee G_{k}$, where $G_{i}$’s are chordal C-I cographs. By Remark 3, there exists a pairwise disjoint set $C_{1}^{i},C_{2}^{i}$ and $C_{3}^{i}$ such that $V(G_{i})=C_{1}^{i}\cup C_{2}^{i}\cup C_{3}^{i}$ and $xy\in E(G_{i})$ if and only if $x,y\in C_{1}^{i}\cup C_{2}^{i}$ or $x,y\in C_{2}^{i}\cup C_{3}^{i}$. Thus in $G$, $\bigcup\limits_{i=1}^{k}C_{2}^{i}$ are universal vertices and for $u,v\in V(G)$ such that $uv\notin E(G)$ if and only if both $u$ and $v$ in $V(G_{i})$ with $u\in C_{1}^{i}$ and $v\in C_{3}^{i}$, for some $i$. Thus, for any $u,v,w\in V(G)$ with $uv,uw\notin E(G)$, then $vw\in E(G)$.
    The structure of an arbitrary C-I cograph is depicted in Fig. 4.
    Fig. 4: General structure of a C-I cograph $G$
    Lemma 10.
    Let $G$ be a C-I cograph then there exists a poset $P=P_{1}+P_{2}+\cdots+P_{r}$, where $P_{i}$’s are completely ranked poset of rank 2 such that $G\cong G_{P}$.
    Proof.
    Let $G$ be a C-I cograph. Then $G$ is the join of chordal C-I cographs. That is, let $G=G_{1}\vee G_{2}\vee\dots G_{r}$. Since $G_{i}$’s are chordal cographs, $G_{i}$’s are C-I graphs of the completely ranked poset $P_{i}$’s of rank at most 2. Hence $P=P_{1}+P_{2}+\cdots+P_{r}$. ∎
    Fig. 5: Claw
    Lemma 11.
    If $G$ is a C-I cograph then $G$ is a claw-free graph.
    Proof.
    Let $G$ be a C-I cograph and $G=G_{1}\vee G_{2}\vee\cdots\vee G_{k}$, where $G_{i}$’s are chordal C-I cograph for $i=1,2,\ldots k$. Suppose that $G$ contains an induced claw by the vertices $\{u,v,w,x\}\subseteq V(G)$ with edges $ux,vx,wx\in E(G)$ and $uv,vw,uw\notin E(G)$ (see Fig.7). Since $G=G_{1}\vee G_{2}\vee\cdots\vee G_{k}$, the non-adjacent vertices lie in the same $G_{j}$ for some $j\in\{1,2,\ldots,k\}$. That is, $u,v,w\in V(G_{j})$. Since $G_{j}$ is a chordal C-I cograph, if $uv,uw\notin E(G_{j})$ then $vw\in E(G_{j})$, which is a contradiction. Hence, the C-I cograph is a claw-free graph. ∎
    A cotree is a tree in which the internal nodes are labelled with the numbers 0 and 1. Every cotree $T$ defines a cograph $G$ having the leaves of $T$ as vertices, and in which the subtree rooted at each node of $T$ corresponds to the induced subgraph in $G$ defined by the set of leaves descending from that node. A subtree rooted at a node labelled 0 corresponds to the union of the subgraphs defined by the children of that node and a node labelled 1 corresponds to the join of the subgraphs defined by the children of that node. The cotree satisfies the property that, on every root-to-leaf path, leaves are the vertices of the graph, the labels of the internal nodes alternate between 0 and 1, and every internal node has at least two children. The cotree can be easily obtained from any tree labelled with such $0/1$ $T$ by coalescing all pairs of child-parent nodes in $T$ having the same label or where the parent has only one child. Vertices $x$ and $y$ of $G$ are adjacent in $G$ if and only if their lowest common ancestor(LCA) in the cotree is labelled 1. This representation is unique and every cograph can be represented in this way by a cotree [13]. Since $G$ is a connected graph, the root of the cotree is a 1-node.
    It is known that cographs have a unique tree representation, called a cotree. Using the cotree, it is possible to design very fast polynomial-time algorithms for problems that are intractable for graphs in general. Such problems include chromatic number, clique determination, clustering, minimum weight domination, isomorphism, minimum fill-in, and Hamiltonicity. A linear-time cograph recognition algorithm, such as those in [10, 14, 21], which also builds a cotree, a data structure that fully encodes the cograph.
    The Fig. 6 illustrates a cograph $G$ and an embedding of the corresponding cotree $T_{G}$. The leaves of $T_{G}$ represent the set of vertex $V(G)$, and each internal node signifies the union (0) or the joining (1) operations in the children. The significance of the 0(1) nodes is captured by the fact that $xy\in E(G)$ if and only if $LCA(x,y)$ is a 1 node, as shown in Fig. 6.
    Fig. 6: $G$ be a cograph and $T_{G}$ be the cotree of $G$
    Fig. 7: Cotree for C-I cograph $(\mathcal{T_{C}})$
    The general structure of the cotree of C-I cographs is shown in Fig. 7, and we denote the family of cotree of C-I cograph as $\mathcal{T_{C}}$. The children of the root node are referred to as level 2 nodes, while the children of level 2 nodes are termed level 3 nodes. Let $G$ be a C-I cograph, $G=G_{1}\vee G_{2}\vee\cdots\vee G_{k}$, where $G_{i}$’s are chordal C-I cographs. Then by Remark 4, two vertices $x$ and $y$ are not adjacent in $G$ if and only if $x\in C_{1}^{i}$ and $y\in C_{3}^{i}$ for $i=1,2,\ldots,k$. Therefore, in the cotree, the lowest common ancestor(LCA) of $x$ and $y$ is 0-node. That is, $LCA(x,y)=0$ for $x\in C_{1}^{i}$ and $y\in C_{3}^{i}$. In other cases, LCA is a 1-node. That is, $LCA(x,y)=0$ for $x\notin C_{1}^{i}$ and $y\notin C_{3}^{i}$.
    Lemma 12.
    The cotree of a C-I cograph which is not a complete graph satisfies the following properties:
    • (i)
      The root node (which is a 1-node) of $\mathcal{T_{C}}$ has leaf nodes and 00-nodes as children. the number of 0-nodes (which are children of the root node) is always less than or equal to the number of leaf nodes (which are children of the root node).
    • (ii)
      Every 0-node has exactly two children (both the children are 1-nodes or both the children are leaves nodes or one child is a leaf node and another child is 1-node).
    • (iii)
      The $1$-nodes other than the root node have only leaf nodes as children.
    • (iv)
      The cotree of $G$ belongs to the family of $\mathcal{T_{C}}$.
    Proof.
    Let $G$ be a C-I cograph which is not a complete graph.
    • (i)
      Since $G$ is not a complete C-I graph, it is either a chordal cograph or joins of chordal cographs. Since there are non-adjacent vertices, there exist some 0-nodes as children of the root node. From the structure of the C-I cograph, there exist universal vertices, which should be leaf nodes of the root node. From Theorem 6, $G=G_{1}\vee G_{2}\vee\cdots\vee G_{k}$, where $G_{i}$ are chordal C-I cographs. The leaf nodes of the subtree rooted by each 0-node contain the non-universal vertices of each $G_{i}$. That is, if there are $k$ 0-nodes, then since the universal vertices of each $G_{i}$, are also universal vertices of $G$, and since each $G_{i}$ has at least one universal vertex, the number of 0-nodes of the root node must be at most the number of leaf nodes as a child nodes. Hence (i) follows.
    • (ii)
      Suppose there are 00-nodes having more than two children. Let $0^{\prime}$ be a 00-node having three child nodes. Let $u,v,w$ be the three leaf nodes from the different branches of the subtree with the root node as the $0^{\prime}$-node. Then by the structure of the C-I cograph, we have that $uv,vw,uw\notin E(G)$. Let $x$ be a leaf node of the root of the cotree $\mathcal{T_{C}}$. Then $x$ is adjacent to $u,v$ and $w$. That is, $\{u,v,w,x\}$ induce a claw in $G$ which is a contradiction to Lemma 11. So every 00-node has no more than two children. In a cotree, every node has a minimum of two child nodes. Hence every 00-node has exactly two child nodes as 1-nodes. Hence (ii) follows.
    • (iii)
      Suppose that there is a $1$-node other than the root node which has a 0-node, say $0^{\prime}$. Consider a subtree $T$ of $\mathcal{T_{C}}$ with rooted in $0^{\prime}$. Let $u$ and $v$ be the vertices of $G$ such that the $LCA(u,v)$ is $0^{\prime}$. The vertices $u$ and $v$ exist since the subtree $T$ has exactly two branches. Now consider the subtree $T^{\prime}$ of $\mathcal{T_{C}}$ containing $T$ with root as a 0-node, say $0^{\prime\prime}$ different from $0^{\prime}$. Let $w$ be a vertex of $G$ such that the $LCA(u,v,w)$ is $0^{\prime\prime}$ ( the vertex $w$ exists since there are two children from $0^{\prime\prime}$ of $T^{\prime}$). Clearly, the vertices $u,v,w$ are mutually non-adjacent. Now any vertex $x$ of $G$ which are leaves of the root node 1 of $\mathcal{T_{C}}$ is adjacent to all of $u,v,w$ so that $u,v,w,x$ form an induced claw, a contradiction. Hence (iii) follows.
    • (iv)
      Follows from (i),(ii) and (iii).
    Theorem 7.
    A cograph $G$ is a C-I cograph if and only if the cotree of $G$ belongs to the family $\mathcal{T_{C}}$.
    Proof.
    It is clear that when $G$ is a complete graph if and only if the cotree is the tree with every vertex of $G$ as a child of the root node and which is a part of $\mathcal{T_{C}}$.
    Let $G$ be a C-I cograph which is not a complete graph. Then by Lemma  12, cotree of $G$ is of the form $\mathcal{T_{C}}$
    Conversely, we need to prove that if $\mathcal{T_{C}}$ is the cotree of a cograph $G$ then $G$ is a C-I graph. Consider the set of leaves of the cotree $\mathcal{T_{C}}$ ( Refer Figure 7 ). It follows by the definition of cotree of a cograph $G$ that the sets $C_{i}^{j}$’s are cliques of $G$, for $i=1,3$ and $j=1,2,\ldots,k$ and $\bigcup\limits_{i=1}^{k}C_{2}^{i}=\{w_{1},w_{2},\ldots,w_{m}\}$ are universal vertices. Now $C_{1}^{j}$ is adjacent to every vertex in $G$ except $C_{3}^{j}$ and $C_{3}^{j}$ is adjacent to every vertex in $G$ except $C_{1}^{j}$ for $j=1,2,\ldots,k$. Then by Remark 4, $G$ is a C-I cograph.
    In [14], Corneil et al. presented a linear-time algorithm for recognizing cographs and constructing their cotree representation. Using that algorithm, we get the cotree.
    Theorem 8.
    [14] The family of cographs possesses a linear-time recognition algorithm by the construction of the cotrees in linear-time.
    Input : A connected graph $G$ with $|V(G)|=n$
    Output : $G$ is a C-I cograph or not. If G is a C-I cograph then find a poset $P$ such that $G_{P}\cong G$.
    • Step 0:
      Using the recognition algorithm in Theorem 8 determine the cotree if $G$ is a cograph and go to Step 1. Otherwise, stop and return $G$ is not cograph
    • Step 1:
      Perform BFS from the root node 1 of the cotree.
    • 1Step 2:
      If BFS completes in one step by reaching all the leaf nodes (if all the adjacent nodes of the root node are
    2 leaf nodes), then the graph is a complete graph and it is C-I cograph; stop, and we are done.
    Otherwise, go to Step 3.
  • Step 3(i):
    After Step 2, only 0-nodes are obtained. Then $G$ is not a C-I cograph and Stop. 3Step 3(ii): After Step 2, both 0-nodes and leaf nodes are obtained. Check whether the number of zero nodes is
  • 4 greater than the number of leaf nodes. If yes, then $G$ is not a C-I cograph and stop.
    Otherwise goto Step 4
  • Step 4:
    Continue the searching at level 2. If 00-nodes has more than two children as $1$-nodes then $G$ is not a C-I cograph and stop. Otherwise, go to Step 5. 5Step 5: Continue searching at level 3. In level 3, if any $1$-node has 00-node as a child, then $G$ is not a C-I cograph
  • and stop. If all the children of the $1$-nodes are the leaf node then we are done. The resulting graph is a C-I cograph.
  • 6Step 6:
    Label the leaf nodes of each 0-node  branch in level 2 as $C_{1}^{i}$ and $C_{3}^{i}$. Partition the leaf nodes from
  • the rooted node 1 into the number of 0-nodes. That is, the leaf nodes are partitioned into $r$ sets, $C_{2}^{i}$ for $i=1,2,\ldots,r$, where $r$ denotes the number of 0-nodes at level 2.
  • 7Step 7:
    Construct completely ranked posets $P_{i}$ with rank 0 elements as $C_{1}^{i}$, rank 1 elements as $C_{2}^{i}$, and rank 2 elements as $C_{3}^{i}$ for $i=1,2,3,\ldots,r$.
  • Return $P=P_{1}+P_{2}+\cdots+P_{r}$ as one of the poset of the C-I cograph $G$.
    Algorithm 2 Algorithm for recognizing C-I cograph
    Proof of correctness: The correctness of the Algorithm 2 up to step 5 follows from the structure of cotree stated in Lemma 12 and Theorem 7, and the related facts mentioned above. If after step 5 results in a C-I cograph, then the poset $P=P_{1}+P_{2}+\cdots+P_{r}$ is certification and it follows from Remark 4, Lemma 10 and Theorem 7 that it is one of the completely ranked poset corresponding to the input graph $G$.
    Now the size of the cotree of a C-I cograph can be estimated as follows. Let $G$ be a C-I cograph with $n$ vertices and $G=G_{1}\vee G_{2}\vee\cdots G_{k}$. Then from the structure of the cotree $T$ of $G$, $T$ contains leaf nodes (which are the $n$-vertices of $G$), 0-nodes, and 1-nodes. Then $T$ contains $k$ 0-nodes, and each 0-node has exactly 2 children. Therefore $T$ has at most $1+k+2k+n$ vertices. The worst case for $k$ is when $k=\frac{n}{3}$. Therefore in the worst case, the total number of vertices in $T$ is $31+\frac{n}{3}+\frac{2n}{3}+$ 3, which is of the order of $O(n)$.
    The time complexity of the recognition algorithm for cographs and finding its cotree by Theorem 8 is linear. Now the complexity of Algorithm 2 is linear in the size of the cotree, which is $O(n)$ in the worst case since the complexity of the BFS can be performed on the cotree in linear-time.
    Acknowledgment:
    Arun Anil acknowledges the financial support from the University of Kerala, for providing University Post Doctoral Fellowship (Ac EVII 5576/2023/UOK dated 30/06/2023).